# Pre-Training Representations of Binary Code Using Contrastive Learning

### v3

@@ 4.1 Embedding Analysis
We perform an embedding analysis to understand the changes in binary code embedding after pre-training our model. The visualization presented in Figure 6 serves as an intuitive case study to illustrate the distribution of embeddings generated by our model. This rough visualization is intended to provide an initial, qualitative glimpse into how distinct functions are scattered across the embedding space, reflecting the model’s ability to differentiate various functionalities. It should be noted that this is not the sole evidence of the model’s effectiveness; comprehensive qualitative and quantitative analyses are presented in subsequent sections to substantiate our claims with robust empirical data. As shown in Figure 6, the five models generated different embeddings for the binary code of the AnghaBench test set. Since the original embedding dimension was 768, we used T-SNE [28] to reduce the dimensionality to two for visualization. As a contrast model, we refer to the model trained solely with primary contrastive learning for 10 epochs as ContraBin (PCL).
General embedding analysis The comparison models, including RoBERTa, CodeBERT, and GraphCodeBERT, showed some clustering of binary code embedding in the two-dimensional space. However, according to the dataset statistics, all binary code in the AnghaBench test set is unique, and therefore should not show any clustering. Thus, we can conclude that the binary code could not be correctly embedded into the manifold by the comparison models. On the other hand, ContraBin (PCL) and ContraBin-generated binary code embeddings showed more distinct properties. The embedding generated by ContraBin (PCL) showed only two natural clusters, while there was only one in ContraBin-generated embedding. The visualization of the embedding proves that our model can improve embedding quality.
Case study To investigate how our model projects binary code into the manifold, we selected two binary codes whose paired source codes had similar semantics. Specifically, we used two source codes called process_record_keymap and process_record_user, respectively, as shown in Figure 6. Both codes had similar input sets and functionality. When we used RoBERTa, CodeBERT, and GraphCodeBERT to project their compiled binary code into the manifold and visualize them, we observed that the two programs were separated by a certain distance, as indicated by the circled red and orange dots. In contrast, our model, which included ContraBin (PCL) and ContraBin, was able to project them into much closer positions in the manifold. In comparison with the embedding generated by ContraBin (PCL), ContraBin was able to further reduce the distance between them, demonstrating an improvement in performance in binary code representation learning by utilizing intermediate contrastive learning.
@@ 4.2 RQ1: Binary Code Analysis
TABLE V: Quantitative evaluation of ContraBin and multiple pre-trained methods on binary code summarization of AnghaBench using BLEU-4, GLUE-4, ROUGE-2, METERO and exact match (xMatch).
@@
RoBERTa [13] 24.30 25.55 30.16 4.95
CodeBERT [12] 32.07 (+31.98%) 32.80 (+28.39%) 38.83 (+28.75%) 7.91 (+60.00%)
GraphCodeBERT [9] 32.43 (+33.44%) 32.94 (+28.92) 39.45 (+30.80) 7.91 (+60.00)
ContraBin (PCL) 30.89 (+27.13%) 31.13 (+21.84%) 37.02 (+22.76%) 6.70 (+35.56%)
ContraBin (Ours) 34.36 (+41.39%) 34.82 (+36.28%) 41.20 (+36.60%) 9.34 (+88.89%)
@@ 4.3 RQ2: Binary Code Comprehension
Result for AnghaBench (Summarization) For the binary code summarization task, we use four commonly used metrics, including BLEU-4 [29], GLUE-4 [30], ROUGE-2 [31], and Exact Match (xMatch) to evaluate the performance of all models. As shown in Table V, ContraBin outperforms all other models in all metrics, achieving substantial improvements of 41.39% for BLEU-4, 36.82% for GLUE, 36.6% for ROUGE, and 88.89% for xMatch over the baseline method.
We further demonstrate the effectiveness of ContraBin through a case study shown in Figure 7. In the first case, ContraBin accurately captures the functionality of the program in summarizing binary code, whereas the summarization of other models have diverged to different objectives, such as cpu, file, lock, etc. In the second case, other models encountered issues in summarizing long binary code. For instance, ContraBin (PCL) generates incomplete sentences, while RoBERTa, CodeBERT, and GraphCodeBERT generate wrong summarization and overly general summaries, respectively. This showcases the long-term semantic consistency and robustness of binary code summarization achieved through the use of binary embeddings generated by ContraBin.
@@ 5 Ablation Studies and Model Insights
TABLE VII: Ablation studies of ContraBin on binary functional algorithm classification of POJ104 using mean of average precision (MAP), recall (MAR) and F1 score (MAF1).
@@
Comment 42.76 (-2.32%) 33.23 (-1.89%) 37.40 (-2.07%)
Anchor 40.96 (-6.44%) 31.07 (-8.27%) 35.34 (-7.46%)
Interpolations 37.03 (-15.42%) 26.99 (-20.31%) 31.22 (-18.25%)
@@ 5.1 Ablation Studies
We assess each component of our model through ablation studies. Specifically, we remove three conditions to test the efficiency of each design element: (1) We remove comments as one of the modalities during training, (2) we replace the anchored encoder with a learnable one so that it can be updated during training, and (3) we eliminate both linear and non-linear interpolation from our model design. For the pretraining process, we adhere to the original hyperparameters, training the first and second ablations for 30 epochs and the third ablation for 10 epochs. For the third ablation, we ensure that the model is converged by the end of pretraining. We test the pretrained models on the binary functional algorithm classification of POJ140, using the same hyperparameters. The results are shown in Figure VII.
From the figure, it is evident that removing comments, stopping gradients, and eliminating interpolation decrease the MAF1 scores by 2.07%, 7.46%, and 18.25%, respectively. This demonstrates the effectiveness of each component in our model design. The baseline performance observed in these ablation studies is as follows: MAP of 43.78, MAR of 33.87, and MAF1 of 38.19. We will use these baseline performance metrics as a reference in the following sub-analysis to further dissect the contributions of individual components.
@@ 5.2 Comparison of Human-Written and LLM-Generated Comments in Pretraining
Performance comparison The experiments reveal a striking decrease in performance when ContraBin is trained with human-written comments compared to LLM-generated comments. Specifically, the use of human-written comments led to a reduction of 40.67% in MAP, 30.86% in MAR, and 35.13% in MAF1. This dramatic decline underscores the effectiveness of LLM-generated comments in enhancing ContraBin’s ability to comprehend and analyze binary code, highlighting the limitations of human-written comments in this context.
@@ 5.5 Extensibility to Contemporary LLMs
The evolving landscape of LLMs necessitates the adaptation of existing methodologies to newer, more advanced architectures. As LLMs like T5 [raffel2020exploring], CodeT5 [25] and its derivatives become increasingly prominent, it is crucial to ensure that techniques initially developed for models like RoBERTa can be effectively extended to these contemporary frameworks. This section outlines the key adaptations required to transition from RoBERTa’s encoder-only architecture to T5’s encoder-decoder structure.
@@ 6 Related Work
In this section, we highlight key areas of related work, including traditional program analysis and neural models for program analysis related to binary code. Furthermore, we also introduce the background and related work that inspire the design of ContraBin, namely large-scale pre-trained embeddings, simplex interpolation, contrastive learning and graph representation learning.
@@
Neural code models In recent years, neural code models have attracted the attention of researchers in software engineering and security. With the assistance of AI [33], it can improve many source code analysis tasks, such as neural clone detection, neural code similarity comparison, neural code search, etc [34, 35, 36, 37, 38]. For example, SySeVR combines syntax, semantics, and neural vector representation to detect software vulnerabilities [39], InnerEye adopts techniques in Natural Language Processing (NLP) to compare between samples of binary code [40], GNN-BPE [41] applies Graph Neural Networks (GNN) on binary code by fusing the semantics of Control Flow Graphs (CFG), Data Flow Graphs (DFG), and call graphs, and Bin2Vec [42] learns binary code representation via Graph Convolutional Networks (GCN). However, these methods involve complex aggregation of additional binary code representations and have limited generalizability across tasks.

### v4

@@ 4.1 Embedding Analysis
We present a case study to analyze the binary code embeddings generated by our model. While binary codes in the AnghaBench test set are unique, semantically related codes can naturally cluster together in the embedding space due to shared functionality or structural patterns. This clustering is an expected property of a well-trained model and demonstrates its ability to encode meaningful relationships among binary codes. This analysis provides an initial qualitative evaluation of the embedding behavior, offering insights into how our model encodes these relationships. While this analysis highlights trends in embedding quality, it is not intended to serve as definitive evidence, and we provide more robust quantitative evaluations in subsequent sections.
Case study: Semantic alignment of binary codes To evaluate the semantic alignment, we selected two binary functions derived from process_record_keymap and process_record_user, which share similar input structures and functionality. As shown in Figure 6, embeddings generated by RoBERTa, CodeBERT, and GraphCodeBERT project these functions far apart in the embedding space. This separation indicates that these models struggle to capture the shared semantic characteristics of the two functions. In contrast, embeddings generated by ContraBin (PCL) position the two functions closer together, reflecting a better alignment of their shared semantics. ContraBin further reduces the distance between these functions, demonstrating its ability to refine binary code embeddings through intermediate contrastive learning.
General trends in embedding space structure To analyze how different models organize binary codes in the embedding space, we examined patterns across the AnghaBench test set. Figure 6 illustrates that embeddings generated by RoBERTa, CodeBERT, and GraphCodeBERT exhibit scattered distributions, with no discernible grouping for binary codes derived from similar tasks, such as hash table operations or data processing routines. In contrast, embeddings from ContraBin (PCL) display a more structured distribution, where binary codes related to similar functionalities begin to cluster into distinct regions. Embeddings generated by ContraBin refine this structure further, consolidating these clusters into a cohesive region while maintaining separation for unrelated binary codes. This visualization reveals distinct patterns in how the embedding spaces produced by different models are structured, offering insights into the organization of binary code representations.
@@ 4.2 RQ1: Binary Code Analysis
TABLE V: Quantitative evaluation of ContraBin and multiple pre-trained methods on binary code summarization of AnghaBench using BLEU-4, GLUE-4, ROUGE-2, and exact match (xMatch). Performance differences are calculated relative to GraphCodeBERT.
@@
RoBERTa [13] 24.30 (-25.05%) 25.55 (-22.45%) 30.16 (-23.52%) 4.95 (-37.41%)
CodeBERT [12] 32.07 (-1.11%) 32.80 (-0.43%) 38.83 (-1.57%) 7.91 (0.00%)
GraphCodeBERT [9] 32.43 (Base) 32.94 (Base) 39.45 (Base) 7.91 (Base)
ContraBin (PCL) 30.89 (-4.75%) 31.13 (-5.50%) 37.02 (-6.15%) 6.70 (-15.28%)
ContraBin (Ours) 34.36 (+5.95%) 34.82 (+5.71%) 41.20 (+4.43%) 9.34 (+18.10%)
@@ 4.3 RQ2: Binary Code Comprehension
Result for AnghaBench (Summarization) For the binary code summarization task, we evaluate all models using four widely adopted metrics: BLEU-4 [27], GLUE-4 [28], ROUGE-2 [29], and Exact Match (xMatch). As shown in Table V, ContraBin achieves consistent improvements across all metrics compared to the baseline method, GraphCodeBERT. Specifically, ContraBin shows improvements of 5.95% in BLEU-4, 5.71% in GLUE-4, 4.43% in ROUGE-2, and 18.08% in xMatch over GraphCodeBERT, showcasing its superior capability in binary code summarization.
To further highlight the effectiveness of ContraBin, we present a case study in Figure 7. In the first example, ContraBin successfully captures the functionality of the binary program, while other models diverge to irrelevant objectives, such as ”cpu,” ”file,” or ”lock.” In the second example, other models struggle with summarizing longer binary code. For instance, ContraBin (PCL) generates incomplete sentences, whereas RoBERTa, CodeBERT, and GraphCodeBERT produce either incorrect or overly general summaries. These cases demonstrate ContraBin’s ability to maintain long-term semantic consistency and robustness in binary code summarization, enabled by its novel binary embeddings.
@@ 5 Ablation Studies and Model Insights
TABLE VII: Ablation studies of ContraBin on binary functional algorithm classification of POJ104 using mean of average precision (MAP), recall (MAR), and F1 score (MAF1). Differences are noted in points (pt).
@@
Human-written comments 3.11 (-40.67 pt) 3.01 (-30.86 pt) 3.06 (-35.13 pt)
Comment 42.76 (-1.02 pt) 33.23 (-0.64 pt) 37.40 (-0.79 pt)
Anchor 40.96 (-2.82 pt) 31.07 (-2.80 pt) 35.34 (-2.85 pt)
Interpolations 37.03 (-6.75 pt) 26.99 (-6.88 pt) 31.22 (-6.97 pt)
@@ 5.1 Ablation Studies
We assess each component of our model through ablation studies. Specifically, we remove three conditions to test the efficiency of each design element: (1) We remove comments as one of the modalities during training, (2) we replace the anchored encoder with a learnable one so that it can be updated during training, and (3) we eliminate both linear and non-linear interpolation from our model design. For the pretraining process, we adhere to the original hyperparameters, training the first and second ablations for 30 epochs and the third ablation for 10 epochs. For the third ablation, we ensure that the model is converged by the end of pretraining. We test the pretrained models on the binary functional algorithm classification of POJ140, using the same hyperparameters. The results are shown in Table VII.
From the table, it is evident that removing comments, stopping gradients, and eliminating interpolation decrease the MAF1 scores by 2.07 pt, 7.46 pt, and 18.25 pt, respectively. This demonstrates the effectiveness of each component in our model design. The baseline performance observed in these ablation studies is as follows: MAP of 43.78, MAR of 33.87, and MAF1 of 38.19. We will use these baseline performance metrics as a reference in the following sub-analysis to further dissect the contributions of individual components.
@@ 5.2 Comparison of Human-Written and LLM-Generated Comments in Pretraining
Performance comparison The experiments reveal a striking decrease in performance when ContraBin is trained with human-written comments compared to LLM-generated comments. Specifically, Table VII shows that the use of human-written comments results in MAP dropping from 43.78 to 3.11, MAR from 33.87 to 3.01, and MAF1 from 38.19 to 3.06. These reductions of 40.67 pt, 30.86 pt, and 35.13 pt, respectively, highlight the stark contrast in effectiveness between the two types of comments.
The collapse in performance with human-written comments underscores the critical role of concise and consistent LLM-generated comments in enhancing ContraBin’s ability to comprehend and analyze binary code. Human-written comments often introduce noise or inconsistencies that hinder representation learning, while the structured nature of LLM-generated comments provides a stronger foundation for binary code analysis. The performance gap can be attributed to the following factors:
  • Redundancy and Verbosity: Human-written comments often include excessive or redundant details, introducing noise that disrupts semantic alignment between source code, comments, and binary code.
  • Focus on Implementation Details: Unlike LLM-generated comments, human-written comments tend to emphasize specific implementation nuances rather than summarizing the core functionality, which impairs generalizable learning.
  • Variability in Style and Structure: Human-written comments exhibit inconsistent styles and structures, creating additional challenges for coherent learning compared to the uniformity of LLM-generated comments.
These findings emphasize the necessity of concise, high-level comments for effective representation learning and highlight the critical impact of training data quality on model performance. Future work should further investigate the interplay between comment quality and model effectiveness to better understand these dynamics.
@@ 5.5 Extensibility to Contemporary LLMs
The evolving landscape of LLMs necessitates the adaptation of existing methodologies to newer, more advanced architectures. As LLMs like T5 [30], CodeT5 [24] and its derivatives become increasingly prominent, it is crucial to ensure that techniques initially developed for models like RoBERTa can be effectively extended to these contemporary frameworks. This section outlines the key adaptations required to transition from RoBERTa’s encoder-only architecture to T5’s encoder-decoder structure.
@@ 6 Related Work
In this section, we highlight key areas of related work, including traditional program analysis techniques and neural models for binary code analysis. Additionally, we discuss the foundational concepts that inspired the design of ContraBin, such as large-scale pre-trained embeddings, simplex interpolation, contrastive learning, and graph representation learning. These areas provide a strong basis for understanding how ContraBin leverages advancements in program analysis and machine learning to improve binary code representation and comprehension.
@@
Neural code models In recent years, neural code models have attracted the attention of researchers in software engineering and security. With the assistance of AI [32], it can improve many source code analysis tasks, such as neural clone detection, neural code similarity comparison, neural code search, etc [33, 34, 35, 36, 37]. For example, SySeVR combines syntax, semantics, and neural vector representation to detect software vulnerabilities [38], InnerEye adopts techniques in Natural Language Processing (NLP) to compare between samples of binary code [39], GNN-BPE [40] applies Graph Neural Networks (GNN) on binary code by fusing the semantics of Control Flow Graphs (CFG), Data Flow Graphs (DFG), and call graphs, and Bin2Vec [41] learns binary code representation via Graph Convolutional Networks (GCN). Compiler chain detection has also emerged as a promising application of machine learning in binary analysis, enabling the identification of the toolchains used to generate binary executables and aiding tasks like provenance analysis and malware detection [42, 43]. Additionally, the survey by Marcelli et al. [44] provides a detailed review of how machine learning techniques are solving the binary function similarity problem, which complements the above-discussed approaches. However, these methods involve complex aggregation of additional binary code representations and have limited generalizability across tasks.
@@ 7 Conclusion
Data Availability Statement
The data and code used in this study are publicly available on Zenodo at https://zenodo.org/records/14435423. The repository includes:
  • Preprocessed datasets used for training and evaluation.
  • Complete implementation of the ContraBin framework, including all scripts for data preprocessing, model training, and evaluation.
  • Detailed documentation and configuration files to facilitate reproducibility and ease of use.
Researchers are encouraged to access the repository to replicate our findings or use the provided resources as a foundation for further studies. Any questions or issues regarding the data or code can be directed to the corresponding authors.